OpenStudy (anonymous):

[17(x^4y^3)^8]/[34(x^5y^2)^4] I know the answer's (x^12y^16)/2, I just forget how to get there! Help please!

OpenStudy (anonymous):

34/17=2 and 32-20=12 and 24-8=16. The equation equates to: (x^12y^16)/2